OpenStudy (dtan5457):

if f(g(x)=2x/(x+z), and g(x)=2x/(x+4) find f(x)

OpenStudy (dtan5457):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

is the `x+z` supposed to be `x+4` ?

OpenStudy (dtan5457):

they are both under 2x so no

jimthompson5910 (jim_thompson5910):

then I'm not sure. That z is throwing me off

OpenStudy (dtan5457):

perhaps its an x+2? I mean my teacher's handwriting does look like a Z.

jimthompson5910 (jim_thompson5910):

that sounds possible

OpenStudy (anonymous):

i have an answer that might work, but i think you are supposed to use inverse functions some how

OpenStudy (anonymous):

by trial and error i came up with \[f(x)=\frac{1}{x-\frac{2}{x}}\]] we can see if that works, and also clean it up a bit

OpenStudy (anonymous):

in any case it is wrong, but i think i can fix it

OpenStudy (anonymous):

lets try this one \[f(x)=\frac{1}{\frac{1}{x}-\frac{2}{2x}}\]

OpenStudy (anonymous):

ok one more time without me being so stupid \[f(x)=\frac{1}{\frac{1}{x}-\frac{2}{x+4}}\]

Nnesha (nnesha):

*

OpenStudy (anonymous):

i think that actually works

OpenStudy (dtan5457):

where would a Z come up though? or is that for x+2

OpenStudy (anonymous):

i used 2 with z it is too weird

OpenStudy (dtan5457):

so it would be impossible with z?

OpenStudy (anonymous):

i am wagering it is 2

OpenStudy (anonymous):

i wonder if we can do it with differentiation though, is that part of this class?

OpenStudy (dtan5457):

this is pre calc we only did difference quotients so far

OpenStudy (anonymous):

maybe @Zarkon has a better idea seems like it needs to use inverse functions some how but i am clogged up in any case i am almost certain the (third) answer i wrote above works, just trying to figure how to subtract 2 in the denominator

OpenStudy (anonymous):

take the reciprocal, subtract 2 in the numerator, then take the reciprocal again of course that mess i wrote can be cleaned up without using compound fractions

OpenStudy (dtan5457):

my teacher may have wrote something wrong of some sort its only been like 2 weeks into school i doubt he would have something harder than you can solve

OpenStudy (anonymous):

usually you are told \(f\) and \(f\circ g\) and asked to find \(g\) that you can do by finding the inverse of \(f\) and applying it this is a bit different

OpenStudy (dtan5457):

you are right in my class example we had to find g and that was fairly simple